# Personalization of Large Language Models: A Survey

### v1

@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools capable of performing a wide range of natural language processing (NLP) tasks with remarkable proficiency (e.g., Radford et al., 2018; Devlin et al., 2019; Lewis et al., 2019; Radford et al., 2019; Brown et al., 2020; Raffel et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Groeneveld et al., 2024). Empirically, these models have demonstrated their capability as generalist models, allowing them to perform numerous tasks such as text generation, translation, summarization, and question-answering with decent accuracy. Notably, LLMs can perform effectively in zero-shot or few-shot settings, meaning they can follow human instructions and perform complex tasks with little to no task-specific training data (Bommasani et al., 2021; Liu et al., 2023c). This capability eliminates the need for extensive fine-tuning of their parameters, thereby significantly simplifying human interaction with machines through straightforward input prompts. For instance, users can engage with LLMs in a conversational format, making interactions more intuitive and accessible. Such robust and versatile abilities of LLMs have led to the creation of numerous applications, including general AI assistants (AutoGPT, 2024), copilots (Microsoft, 2024), and personal LLM-based agents (Li et al., 2024f). These applications assist users in a wide range of activities such as writing emails, generating code, drafting reports, and more.
@@
Personalization of LLMs has recently gained a lot of attention (Salemi et al., 2023; Tseng et al., 2024). However, existing research on personalized LLMs generally falls into two categories: (a) personalized text generation, which focuses on producing single or multi-turn text within a personalized context, and (b) the application of LLMs in downstream personalization tasks, such as recommendation systems. A large number of studies in these two areas have evolved independently, often without a cohesive or unified perspective. Additionally, existing surveys (Chen, 2023; Chen et al., 2024b; c) tend to focus exclusively on one of these aspects, lacking a comprehensive, unified perspective that systematically defines the key components and synthesizes insights across both dimensions of personalized LLMs. Although these two areas differ in characteristics and objectives, a unified perspective on personalized LLMs is crucial as it can bridge the gap between these two research communities, promoting collaboration and knowledge sharing that leads to more generalizable and versatile systems. For example, advancements in user-specific text generation can improve recommendation systems by enabling more personalized and explainable suggestions through conversational interactions. By integrating insights from both areas, researchers can develop LLMs that not only generate text tailored to individual preferences but also improve user satisfaction across diverse applications. This interdisciplinary approach fosters more holistic solutions, addressing both personalization and performance in a complementary way. In this work, we unify the literature from these different fields by proposing intuitive taxonomies for the granularity of personalization, techniques, evaluation, datasets, and usage scenarios of personalized LLMs.
@@
    An analysis and taxonomy of the personalization granularity of LLMs (Section 4). We propose three different levels of personalization granularity for LLMs, including (i) user-level personalization, (ii) persona-level personalization, and (iii) global preference personalization. We formalize these levels, and then discuss and characterize the trade-offs between the different granularities of LLM personalization. Notably, user-level personalization is the finest granularity; however, it requires a sufficient amount of user-level data. In contrast, persona-level personalization groups users into personas and tailors the experience based on persona assignments. While it doesn’t provide the same granularity as user-level personalization, it is effective for personalizing experiences for users with limited data. Finally, global preference alignment caters to overall preferences by the general public and does not offer user-specific personalization. [^1] We include it here for completeness, though it is not the focus of this work.
@@
    A survey and taxonomy of techniques for LLM personalization (Section 5). We categorize and provide a comprehensive overview of the current techniques for personalizing LLMs based on how user information is utilized. Our taxonomy covers various categories of methods such as retrieval-augmented generation (RAG), prompt engineering, supervised fine-tuning, embedding learning, and reinforcement learning from human feedback (RLHF). For each method, we discuss their unique characteristics, applications, and the trade-offs involved. Our detailed analysis helps in understanding the strengths and limitations of different personalization techniques and their suitability for various tasks.
@@
Figure 1: Taxonomy for Personalized LLM Usage. To bridge the gap in the existing literature on personalized LLMs, we propose the intuitive taxonomy outlined above, which categorizes work into two main areas. The first focuses on studying the personalized text generated directly, while the second emphasizes using personalized information as intermediate steps or implicitly as embeddings to improve the quality of a downstream task such as recommendation systems. See Section 2 for a detailed discussion.
@@ 2 Unifying Personalized LLMs
To bridge the gap between the two distinct lines of work in the literature, we propose an intuitive taxonomy (see Figure 1) that categorizes personalized LLM efforts into two main categories: personalized text generation, and downstream task personalization. In the first category, personalized text generation, the goal is to generate text that directly aligns with individual or group preferences (Salemi et al., 2023; Kumar et al., 2024). For example, a personalized mental health chatbot should generate empathetic responses based on a user’s previous conversations, adapting the tone and language to reflect their emotional state. The focus is on producing personalized content, which is evaluated by assessing the quality of the generated text itself using human evaluation or text-based metrics which might require ground-truth text written by users for evaluation, as the generated text should match or approximate the style or content the user would produce. In the second category, downstream task personalization, personalized LLMs are used to enhance the performance of a specific task, such as recommendation (Lyu et al., 2023; Bao et al., 2023). For instance, an LLM-enhanced movie recommendation system might suggest new films by analyzing a user’s viewing history, preferences, and interactions with previous recommendations. In this scenario, the LLM may generate intermediate tokens or embeddings that enhance the system’s performance on specific downstream tasks. While these intermediate tokens are not evaluated directly, they serve as crucial steps in improving the overall effectiveness of the task-specific system. The performance is assessed through task-specific metrics like recommendation accuracy or task success rates. Unlike the first category, this line of work focuses on improving task outcomes rather than the text generation process itself.
Direct Personalized Text Generation: While there are many categories of techniques for generating personalized text from LLMs, we use a simple RAG-based framework in this definition. For user $i$, we typically have a set of user documents (user-written text), user attributes that may be static, and user interactions, as depicted in Figure 1. The user provides a textual input $x$, that get passed through a query generation function $\phi_{q}$ to obtain the query used to retrieve the top-$k$ elements from the users’ data via the use of a retrieval model $\mathcal{R}$. The retrieval model $\mathcal{R}$ takes the transformed input $\phi_{q}(x)$, users data $\mathcal{D}_{i}$, and a parameter $k$ indicating the number of top entries to return. More formally,
$\mathcal{D}_{i}^{(k)}=\mathcal{R}(\phi_{q}(x),\mathcal{D}_{i},k)$ (1)
where $\mathcal{D}_{i}^{(k)}$ is the set of top-$k$ most similar data for user $i$ from $\mathcal{D}_{i}$. Note that $\phi_{q}$ represents the query generation function, such as query expansion, introduced here for generality; it can also simply be the identity function, giving $\mathcal{D}_{i}^{(k)}=\mathcal{R}(x,\mathcal{D}_{i},k)$. We also note two other aspects. First, while the retrieval model $\mathcal{R}$ shown in Figure 1 is over only user $i$’s data, in general, it can be over any useful set of data that can improve the personalized text generated for that user. Second, the retrieval model $\mathcal{R}$ can also be used over each of the different data types individually (user documents/written text, user attributes, user interactions, etc), and then combined in a certain way. There may also be different retrieval models used for each data type as needed.
Given the retrieved data $\mathcal{D}_{i}^{(k)}$ for user $i$ along with their input text $x$, we generate a personalized prompt for user $i$ and input $x$ using a personalized prompt generation function $\phi_{p}$ that combines $x$ with $\mathcal{D}_{i}^{(k)}$ the top-$k$ user-specific retrieved data. More formally,
$\bar{x}=\phi_{p}(x,\mathcal{R}(\phi_{q}(x),\mathcal{D}_{i},k))$ (2)
where $\bar{x}$ is the new input personalized for user $i$ based on the initial users input $x$ and the retrieved users data $\mathcal{D}_{i}^{(k)}$.
Now, given the personalized input $\bar{x}$ for user $i$ from Eq. 2, we can leverage an arbitrary LLM $\mathcal{M}$ to obtain the personalized text generated $\hat{y}$ for user $i$ [^2] As an aside, $\bar{x}$ from Eq. 2 can also be used as $(\bar{x},y)$ to fine-tune/train an arbitrary LLM $\mathcal{M}$ as well. Hence, the personalized text generated is:
$\hat{y}$ $=$ $\mathcal{M}(\phi_{p}(x,\mathcal{R}(\phi_{q}(x),\mathcal{D}_{i},k)))$
$=$ $\mathcal{M}(\bar{x})$
Some existing works focus entirely on generating this personalized user-specific text $\hat{y}$ and then evaluate the quality of it using the actual text written by the user denoted as $y$ (Salemi et al., 2023). Such approaches directly use the personalized text generated. Therefore, the personalized text generated $\hat{y}$ is evaluated directly against the known ground-truth text $y$ written by the specific user $i$, which we denote in Figure 1 as $\mathbb{E}(\hat{y},y)$. where $\mathbb{E}$ is an evaluation metric such as ROUGE-1, ROUGE-L, METEOR, or any other future metric specifically designed for the personalized text generation task at hand. In general, the goal of $\mathbb{E}$ is to quantify how well the generated text $\hat{y}$ personalized for user $i$’s input $x$ is captured where $y$ is the actual text written by that user.
@@
While evaluating how well the personalized text generated, $\hat{y}$, matches the known ground-truth text, $y$, is crucial, it remains particularly challenging due to the scarcity of datasets with high-quality, user-written labels. This likely contributes to the limited focus on the foundational task of personalized text generation in the literature. Instead, there have been many more works that focus on utilizing the personalized text generated $\hat{y}$ indirectly to improve a downstream task such as recommendation or prediction in general. It is important to note that in these indirect approaches, where the focus is on improving a downstream task, the personalized text output $\hat{y}$ is typically not evaluated and is considered less critical. The key consideration is that the generated intermediate text or its embedding when applied to a downstream task, can enhance the overall system’s performance. While this line of work lacks interpretability regarding the intermediate information generated by LLMs, it has been demonstrated that augmenting systems with this information typically improves performance in downstream personalization-related applications. For a detailed discussion of the evaluation specific to personalized LLMs, see Section 6.
@@ 3.1 Preliminaries
Let $\mathcal{M}$ be an LLM parameterized by $\theta$, which takes a text sequence $X=(x_{1},\cdots,x_{m})\in\mathbb{X}$ as input and produces an output $\hat{Y}\in\hat{\mathbb{Y}}$, where $\hat{Y}=\mathcal{M}(X;\theta)$. The form of $\hat{Y}$ depends on the specific task, with $\hat{\mathbb{Y}}$ representing the output space of possible generations. The inputs can be drawn from a labeled dataset $\mathcal{D}={(X^{(1)},Y^{(1)}),\cdots,(X^{(N)},Y^{(N)})}$, or from an unlabeled dataset of prompts for sentence continuations or completions $\mathcal{D}={X^{(1)},\cdots,X^{(N)}}$. For this and other notation, see Table 2.
@@
A large language model (LLM) $\mathcal{M}$, parameterized by $\theta$, is a multi-layer Transformer model with hundreds of billions (or more) parameters. It can be structured with an encoder-only, decoder-only, or encoder-decoder architecture and is trained on extensive corpora comprising a vast number of natural language tokens (Zhao et al., 2023; Gallegos et al., 2024).
@@
Downstream Tasks $\mathcal{F}$ are specific applications or tasks that utilize the outputs of a model, such as an LLM, to perform practical, real-world functions. These tasks can include but are not limited to classification, translation, recommendation, and information retrieval. Given an input $X$ and the model-generated output $\hat{y}$, a downstream task $\mathcal{F}$ evaluates or utilizes $\hat{y}$ to produce a final prediction $\hat{r}$:
@@
After pre-training with large-scale unlabeled corpora in an unsupervised manner, the resulting in-context-aware word representations are very effective as general-purpose semantic features for a wide range of NLP tasks. With the scaling of their size and techniques such as instruction tuning (Ouyang et al., 2022; Zhang et al., 2023c; Longpre et al., 2023; Zhou et al., 2024a) and RLHF (Christiano et al., 2017; Stiennon et al., 2020b; Rafailov et al., 2024), they exhibit many emergent abilities (Wei et al., 2022a). This enables LLMs to solve complex tasks and engage in natural conversations with humans, even in a zero-shot manner through text prompting for a wide range of downstream tasks such as sequence classification, text generation, and recommendation (Qin et al., 2023). To further enhance LLMs’ performance on specific downstream tasks, models are often fine-tuned with a relatively small amount of task-specific data following the “pre-train, then fine-tune” paradigm, which generally adapts LLMs to particular tasks and achieves better results (Bommasani et al., 2021; Min et al., 2023; Liu et al., 2023b).
@@
A Prompt $\mathcal{H}$ is a specific input or set of instructions provided to a language model, which guides its generation of text $M_{\theta}(\mathcal{H})$. Prompts can vary in complexity from simple word or phrase completions to detailed, structured contexts or questions aimed at eliciting specific types of responses or performing certain tasks. Prompts can be multi-modal, including text, image, audio, or video inputs.
@@
Definition 4 (Recommendation System).
A Recommendation System $RecSys$ is an information filtering system that addresses the problem of information overload by filtering vital information fragments from large amounts of dynamically generated information according to a user’s preferences, interests, or observed behavior about items. For a specific user $i$, given the user-item interaction history $I_{i}$, the user’s profile attributes $a_{i}$, and the user-written text $t_{i}$, a recommendation system aims to predict the user’s preference for a new item (Bobadilla et al., 2013).
@@ 3.2 Formulation of Personalization
User Preferences refer to the specific likes, dislikes, interests, and priorities of an individual user or a group of users. These preferences guide the personalization process by informing the system about the desired characteristics and features of the output. In the context of LLMs, user preferences can be derived from explicit feedback, historical interactions, and contextual signals to tailor responses and improve the relevance and satisfaction of the generated content.
@@
Personalization, the practice of tailoring experiences to the preferences of individual users or groups of users, is crucial for bridging the gap between humans and machines (Rossi et al., 1996; Montgomery et al., 2004; Chen et al., 2024c). Such experiences can include aligning with specific user or group preferences, adjusting the style or tone of generated content, and providing recommendation items based on the user’s interaction history in a wide range of downstream tasks. Users can be actual individuals with a history of interactions or described by specific characteristics such as demographic information, allowing both humans and machines to better understand and cater to their needs. In this work, instead of just focusing on personalization for single individual users, we aim to formalize and clarify the term “personalization” by categorizing its objectives based on the size of the targeted group. We classify personalization into three categories based on their focus: aligning with the preferences of individual users, groups of users, or the general public (Section 4). Additionally, these three levels of personalization enable the incorporation of different types of input data, each contributing uniquely to the personalization process. It is important to note that not all fine-tuning equates to personalization. For example, most supervised fine-tuning practice is a process where models are trained on specific datasets to perform better on a downstream task. However, only fine-tuning that adjusts a model to cater to specific user or group preferences—such as adapting a model to a user’s writing style or content preferences—counts as personalization. In contrast, fine-tuning on a general corpus to improve overall task performance is not personalized, as it does not address the unique preferences of individuals or groups. This distinction is key to understanding the objectives of personalized LLMs across the different levels of granularity.

### v3

@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools capable of performing a wide range of natural language processing (NLP) tasks with remarkable proficiency (e.g., Radford et al., 2018; Devlin et al., 2019; Lewis et al., 2019; Radford et al., 2019; Brown et al., 2020; Raffel et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Groeneveld et al., 2024). Empirically, these models have demonstrated their capability as generalist models, allowing them to perform numerous tasks such as text generation, translation, summarization, and question-answering with decent accuracy. Notably, LLMs can perform effectively in zero-shot or few-shot settings, meaning they can follow human instructions and perform complex tasks with little to no task-specific training data (Bommasani et al., 2021; Liu et al., 2023c). This capability eliminates the need for extensive fine-tuning of their parameters, thereby significantly simplifying human interaction with machines through straightforward input prompts. For instance, users can engage with LLMs in a conversational format, making interactions more intuitive and accessible. Such robust and versatile abilities of LLMs have led to the creation of numerous applications, including general AI assistants (AutoGPT, 2024), copilots (Microsoft, 2024), and personal LLM-based agents (Li et al., 2024h). These applications assist users in a wide range of activities such as writing emails, generating code, drafting reports, and more.
@@
Personalization of LLMs has recently attracted significant attention (Salemi et al., 2023; Tseng et al., 2024), with research primarily focusing on two directions: (a) personalized text generation, which tailors generated text to user-specific contexts, and (b) downstream task personalization, which leverages LLM capabilities to improve performance on targeted applications such as recommendation systems. Despite the extensive research efforts, these two areas have historically developed independently due to technical limitations and methodological differences, often resulting in existing surveys (Chen, 2023; Chen et al., 2024b; c) examining each aspect in isolation. However, these two domains are not fundamentally distinct, as LLMs possess the flexibility to adapt to a wide range of tasks (Qin et al., 2023). Rather, as LLM capabilities continue to evolve, we envision that those two directions will increasingly converge, enabling unified systems in which a single intelligent agent seamlessly transitions from engaging in personalized conversations to reasoning over structured knowledge like product catalogs for task-oriented recommendations. Bridging this current conceptual gap by synthesizing insights across both dimensions thus constitutes an essential step toward creating fully integrated, adaptable, and generalizable user experiences.
To fully understand LLM personalization, it is important to examine these research directions within a unified framework that captures the broader landscape of personalization. Beyond integrating different approaches, a more in-depth discussion is needed on the foundational concepts, techniques, datasets, and evaluation methods that support LLM personalization. Additionally, real-world challenges such as balancing personalization with privacy concerns, mitigating biases, and handling data limitations must be addressed to ensure practical and ethical deployment. In this survey, we provide a comprehensive perspective by introducing systematic taxonomies that categorize personalization based on granularity, methodology, and evaluation strategies. We also explore open problems and potential research opportunities. Through this framework, we connect the various aspects of personalization and offer a structured reference that outlines the essential components needed to develop and evaluate personalized LLMs.
@@
    An analysis and taxonomy of the personalization granularity of LLMs (Section 4). We propose three different levels of personalization granularity for LLMs, including (i) user-level personalization, (ii) persona-level personalization, and (iii) global preference personalization. We formalize these levels, and then discuss and characterize the trade-offs between the different granularities of LLM personalization. Notably, user-level personalization is the finest granularity; however, it requires a sufficient amount of user-level data. In contrast, persona-level personalization groups users into personas and tailors the experience based on persona assignments. While it doesn’t provide the same granularity as user-level personalization, it is effective for personalizing experiences for users with limited data. Finally, global personalization caters to the overall preferences of the general public and does not offer user-specific personalization. [^1] We include it here for completeness, though it is not the focus of this work.
@@
    A survey and taxonomy of techniques for LLM personalization (Section 5). We categorize and provide a comprehensive overview of the current techniques for personalizing LLMs based on how user information is utilized. Our taxonomy covers various categories of methods such as retrieval-augmented generation (RAG), prompt engineering, supervised fine-tuning, embedding learning, and reinforcement learning from human feedback (RLHF). For each category of methods, we discuss their unique characteristics, applications, and the trade-offs involved. Our detailed analysis helps in understanding the strengths and limitations of different personalization techniques and their suitability for various tasks.
@@
Figure 1: Taxonomy for Personalized LLM Usage. To bridge the gap in the existing literature on personalized LLMs, we propose the intuitive taxonomy outlined above, which categorizes work into two main areas. The first focuses on studying the personalized text generated directly, while the second emphasizes using personalized information as intermediate steps or implicitly as embeddings to improve the quality of a downstream task such as recommendation systems. See Section 2 for a detailed discussion. An example of an adaptation function $A$ is a retrieval module. Please note that $y$ here can represent user-written text if available, or alternatively, user preferences and separate reward models that reflect user judgments.
@@ 2 Unifying Personalized LLMs
To bridge the gap between the two distinct lines of work in the literature, we propose an intuitive taxonomy (see Figure 1) that categorizes personalized LLM efforts into two main categories: personalized text generation, and downstream task personalization. In the first category, personalized text generation, the goal is to generate text that directly aligns with individual or group preferences (Salemi et al., 2023; Kumar et al., 2024). For example, a personalized mental health chatbot should generate empathetic responses based on a user’s previous conversations, adapting the tone and language to reflect their emotional state. The focus is on producing personalized content, which is evaluated by assessing the quality of the generated text itself using user-written text if available, or alternatively, user preferences and separate reward models that reflect user judgments, as the generated text should match or approximate the style or content the user would produce. In the second category, downstream task personalization, personalized LLMs are used to enhance the performance of a specific task, such as recommendation (Lyu et al., 2023; Bao et al., 2023). For instance, an LLM-enhanced movie recommendation system might suggest new films by analyzing a user’s viewing history, preferences, and interactions with previous recommendations. In this scenario, the LLM may generate intermediate tokens or embeddings that enhance the system’s performance on specific downstream tasks. While these intermediate tokens are not evaluated directly, they serve as crucial steps in improving the overall effectiveness of the task-specific system. The performance is assessed through task-specific metrics like recommendation accuracy or task success rates. Unlike the first category, this line of work focuses on improving task outcomes rather than the text generation process.
Direct Personalized Text Generation: While there are various techniques for generating personalized text from LLMs, We introduce a general adaptation function $A$ that integrates user-specific information into personalized text generation. An example of $A$ is a retrieval module, with additional concrete examples provided in Figure 7. For a given user $u$, this information may include user-written documents, static attributes, interaction histories, or preferences, as illustrated in Figure 1. Given a user’s textual input $x$, a query generation function $\phi_{q}$ transforms this input to effectively integrate relevant user-specific data via the adaptation function $A$. The adaptation function $A$ leverages the transformed query $\phi_{q}(x)$, user data $\mathcal{D}_{u}$, and optionally a parameter $k$ for flexibility. Subsequently, another transformation function, the personalized prompt generation function $\phi_{p}$, combines the original input $x$ and the output of the adaptation function to form a personalized input $\bar{x}$. Ultimately, the personalized text $\hat{y}$ generated by an LLM $\mathcal{M}$ is:
$\hat{y}=\mathcal{M}(\phi_{p}(x,\mathcal{A}(\phi_{q}(x),\mathcal{D}_{u},k)))=\mathcal{M}(\bar{x})$ (1)
Some existing approaches focus exclusively on generating this personalized text $\hat{y}$ and then directly evaluating its quality. Evaluation typically involves comparing $\hat{y}$ to some form of ground truth or user-specific reference, denoted as $y$. The reference $y$ may represent actual user-written text when available, or more generally, it may encompass user preferences or evaluations obtained from dedicated reward models or user judgments. The evaluation is performed using a metric $\mathbb{E}(\hat{y},y)$, such as ROUGE-1, ROUGE-L, METEOR, or other specialized metrics designed for personalized text generation tasks.
@@
While evaluating how well the personalized text generated, $\hat{y}$, matches the known user preferences is crucial, it remains particularly challenging due to the scarcity of datasets with high-quality, user-written labels. This likely contributes to the limited focus on the foundational task of personalized text generation in the literature. Instead, there have been many more works that focus on utilizing the personalized text generated $\hat{y}$ indirectly to improve a downstream task such as recommendation or prediction in general. It is important to note that in these indirect approaches, where the focus is on improving a downstream task, the personalized text output $\hat{y}$ is typically not evaluated and is considered less critical. The key consideration is that the generated intermediate text or its embedding when applied to a downstream task, can enhance the overall system’s performance. While this line of work lacks interpretability regarding the intermediate information generated by LLMs, it has been demonstrated that augmenting systems with this information typically improves performance in downstream personalization-related applications. For a detailed discussion of the evaluation specific to personalized LLMs, see Sec. 6.
@@ 3.1 Preliminaries
Let $\mathcal{M}$ be an LLM parameterized by $\theta$, which takes a text sequence $X\in\mathbb{X}$ as input and produces an output sequence $\hat{Y}\in\hat{\mathbb{Y}}$, where $\hat{Y}=\mathcal{M}(X;\theta)$. The form of $\hat{Y}$ depends on the specific task, with $\hat{\mathbb{Y}}$ representing the output space of possible generations. The inputs can be drawn from a labeled dataset $\mathcal{D}={(X^{(1)},Y^{(1)}),\cdots,(X^{(N)},Y^{(N)})}$, or from an unlabeled dataset of prompts for sentence continuations or completions $\mathcal{D}={X^{(1)},\cdots,X^{(N)}}$. For this and other notation, see Table 2.
@@
A large language model (LLM) $\mathcal{M}$, parameterized by $\theta$, is a multi-layer Transformer model with billions (or more) of parameters. It can be structured with an encoder-only, decoder-only, or encoder-decoder architecture and is trained on extensive corpora comprising a vast number of natural language tokens (Zhao et al., 2023; Gallegos et al., 2024).
@@
A downstream task is a specific practical application or goal, such as classification, translation, recommendation, or information retrieval, that uses outputs generated by a model (e.g., an LLM). Formally, for a downstream task, we define a corresponding downstream model or function $\mathcal{F}$ that takes as input the model’s output $\hat{y}$ (generated from an initial input $X$) and produces a final result or prediction
@@
After pre-training with large-scale unlabeled corpora in an unsupervised manner, the resulting in-context-aware word representations are very effective as general-purpose semantic features for a wide range of NLP tasks. With the scaling of their size and techniques such as instruction tuning (Ouyang et al., 2022; Zhang et al., 2023c; Longpre et al., 2023; Zhou et al., 2024a) and RLHF (Christiano et al., 2017; Stiennon et al., 2020b; Rafailov et al., 2024), LLMs exhibit many emergent abilities (Wei et al., 2022a). This enables LLMs to solve complex tasks and engage in natural conversations with humans, even in a zero-shot manner through text prompting for a wide range of downstream tasks such as sequence classification, text generation, and recommendation (Qin et al., 2023). To further enhance LLMs’ performance on specific downstream tasks, models are often fine-tuned with a relatively small amount of task-specific data following the “pre-train, then fine-tune” paradigm, which generally adapts LLMs to particular tasks and achieves better results (Bommasani et al., 2021; Min et al., 2023; Liu et al., 2023b).
@@
A Prompt $\mathcal{H}$ is a specific input or set of instructions provided to a language model, which guides its generation of text $M(X;\theta)$. Prompts can vary in complexity from simple word or phrase completions to detailed, structured contexts or questions aimed at eliciting specific types of responses or performing certain tasks. Prompts can be multi-modal, including text, image, audio, or video inputs.
@@ 3.2 Formulation of Personalization
User Preferences refer to the specific likes, dislikes, interests, and priorities of an individual user or a group of users. These preferences guide the personalization process by informing the system about the desired characteristics and features of the output. In the context of LLMs, user preferences can be derived from explicit feedback (e.g., pairwise comparison), historical interactions, and contextual signals to tailor responses and improve the relevance and satisfaction of the generated content.
@@
Personalization, the practice of tailoring experiences to the preferences of individual users or groups of users, is crucial for bridging the gap between humans and machines (Rossi et al., 1996; Montgomery et al., 2004; Chen et al., 2024c). Such experiences can include aligning with specific user or group preferences, adjusting the style or tone of generated content, and providing recommendation items based on the user’s interaction history in a wide range of downstream tasks. Users can be actual individuals with a history of interactions or described by specific characteristics such as demographic information, allowing both humans and machines to better understand and cater to their needs. In this work, instead of just focusing on personalization for single individual users, we aim to formalize and clarify the term “personalization” by categorizing its objectives based on the size of the targeted group. We classify personalization into three categories based on their focus: aligning with the preferences of individual users, groups of users, or the general public (Sec. 4). Additionally, these three levels of personalization enable the incorporation of different types of input data, each contributing uniquely to the personalization process. It is important to note that not all fine-tuning equates to personalization. For example, most supervised fine-tuning practice is a process where models are trained on specific datasets to perform better on a downstream task. However, only fine-tuning that adjusts a model to cater to specific user or group preferences—such as adapting a model to a user’s writing style or content preferences—counts as personalization. In contrast, fine-tuning on a general corpus to improve overall task performance is not personalized, as it does not address the unique preferences of individuals or groups. This distinction is key to understanding the objectives of personalized LLMs across the different levels of granularity.